Please help! I need a step by step solution

2. A dose of 250 mg in 6 mL has been ordered diluted to 50 mL and infused in 45 minutes.

A 15 mL flush follows.

Medication is placed in a burette.

Microdrop tubing is used.

(Do not include the flush in your calculations)


The correct rate is _____gtt/min

And _____mL/hr.
Answer #1

Drug= 250mg in 6ml

Dilution= 50ml

Time : 45minutes or 0.75 hour

Drop factor: micro drop(60)

Formula= amount of fluid× drop factors/ time× minutes

= 50× 60/ 0.75×60

= 66.6 drops per minute.

So 66.6 gtt/ min

1ml equal to 20 drops

So,

66.6/20= 3.3ml/ min

One hour= 60min

One hour= 60× 3.3ml

= 200ml/ hour

Answer:

66.6 gtt/ min

200ml/ hour